OpenStudy (anonymous):

Which is an equivalent expression to positive eighteen minus negative twenty-nine? negative eighteen plus positive twenty-nine positive twenty-nine plus positive eighteen negative eighteen minus positive twenty-nine negative twenty-nine minus negative eighteen

OpenStudy (anonymous):

please somebody

OpenStudy (anonymous):

They all equal 47 ughhh

OpenStudy (anonymous):

@AriPotta can you help

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

Also will fan

OpenStudy (anonymous):

@recon14193 i will medal and fan if you answer but do you understand

OpenStudy (anonymous):

so I think it might be a good Idea to use the numbers rather than type it out first

OpenStudy (anonymous):

18-(-29) this is equal to 18+29=47

OpenStudy (anonymous):

so that is your first indicator of what could be an equivalent expression but as a double check we can look at the other answers' solutions

OpenStudy (anonymous):

-18+29=11

OpenStudy (anonymous):

we already saw how 18+29 would be equivalent

OpenStudy (anonymous):

-18-(+29) the plus is irrelevant so we can write it as -18-29=-47

OpenStudy (anonymous):

-29-(-18)=-29+18=-11

OpenStudy (anonymous):

so we can see the only one equal here would be 29+18

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

this is something you can verify with the calculator but I would recommend in the future writing out your work

OpenStudy (anonymous):

do you understand where you made your errors though

OpenStudy (anonymous):

Yes i do

OpenStudy (anonymous):

alright good